OpenStudy (anonymous):

Solve. x2 + 5x + 6 = 0 A. {1, –6} B. {–1, 6} C. {2, 3} D. {–2, –3}

OpenStudy (zzr0ck3r):

plug the numbers in, or use the ac method.

OpenStudy (zzr0ck3r):

1*6=6 = 2*3 2+3 = 5 so \(x^2+5x+6=0\implies x^2+2x+3x+6=0 \implies x(x+2)+3(x+2)=0\) can you finish?